OpenStudy (anonymous):

What is the volume of a sphere with a 15 ft radius?

OpenStudy (bibby):

formula for a sphere's volume = \(\frac{4}{3} \pi r^3\)

OpenStudy (anonymous):

plug in r=15 in the formula

OpenStudy (anonymous):

and chug on

OpenStudy (jdoe0001):

plug and chug be happy, eat ice-cream, have a brain-freeze then relax

OpenStudy (anonymous):

^dope

OpenStudy (bibby):

step 4 do dope or lsd expand your miiiind

OpenStudy (anonymous):

im still confuzzled can you put it in an easier context

OpenStudy (bibby):

V=\(\frac{4}{3} \pi \text{radius}^3\)

OpenStudy (anonymous):

so 4 divided by 3 x radius cubed?

OpenStudy (bibby):

exactly. what is the radius cubed?

OpenStudy (anonymous):

it would be 3375

OpenStudy (anonymous):

then what do I do

OpenStudy (anonymous):

i'm not a genius so u gotta take it step by step lol

OpenStudy (bibby):

so we plug it in 4/3*pi*3375

OpenStudy (anonymous):

1413 sumthin

OpenStudy (bibby):

get a calculator yo

OpenStudy (bibby):

\(\large \frac{4}{3}* \pi * 3375\)=4500\(\pi\)

OpenStudy (anonymous):

lol i kno diss haha